Evaluate the integral, I  (4x  2y) dxdy 3 in the region R bounded by x  0, x  2, y  x and y  x  2 .


1
Expert's answer
2021-07-11T17:31:59-0400
"\displaystyle\int_{0}^{2}\displaystyle\int_{x}^{x^2}4x^2ydydx=\displaystyle\int_{0}^{2}[2x^2y^2]\begin{matrix}
 x^2 \\
 x
\end{matrix}dx"

"=2\displaystyle\int_{0}^{2}(x^6-x^4)dx=2\big[\dfrac{x^7}{7}-\dfrac{x^5}{5}\big]\begin{matrix}
 2 \\
 0
\end{matrix}"

"=2(\dfrac{2^7}{7}-\dfrac{2^5}{5})=\dfrac{64}{35}(20-7)=\dfrac{832}{35}"